if you apply a froce of 2,200 N to do 175J of work on a car, how far did you move it?
Answer by Gogonati(855)

As you know W=FxD, where, W is the work, F the force and D the distance.
From this formula we find D=W/F, substitute and have: D=175J/2200N=0.008 m.
Answer:The car will move 8 millimeter.